Calculate the electric field in a copper wire of cross-sectional area 2.0 mm^2 carrying a current of 1 A. The resistivity of copper = 1.7 × 10^8 Ω-m.

Grade:upto college level

1 Answers

Aditi Chauhan
9 years ago
Sol. A = 2 * 10^-6 m^2, l = 1 A f = 1.7 * 10^-8 Ω-m E = ? R = fℓ/A = 1.7 * 10^-8 * ℓ/2 * 10^-6 V = IR = 1 * 1.7 * 10^-8 * ℓ/2 * 10^-6 E = dV/dL = V/l = 1.7 * 10^-8 * ℓ/2 * 10^-6 ℓ = 1.7/2 * 10^-2V/m = 8.5 mV/m.